OpenStudy (anonymous):

250x^3-128y^3 factor

OpenStudy (ybarrap):

We can use the difference of cubes formula: \(a^3-b^3=(a-b)(a^2+ab+b^2)\). Here \(a = 125^{1/3}x\) and \(b = 64^{1/3}y\): $$ 250x^3-128y^3\\ =2(125x^3-64y^3)\\ \text{Applying formula above:}\\ =2(125^{1/3}x-64^{1/3}y)(125^{2/3}x^2+125^{1/3}x64^{1/3}y+64^{2/3}y^2)\\ =2(5x-4y)(5^2x^2+5x4y+4^2y^2)\\ =2(5x-4y)(25x^2+20xy+16y^2)\\ $$ That's it! Let me know if you have any questions.